OpenStudy (anonymous):

Suppose the point (4,-3) is on the graph y=f(x). Find a point on graph y= 6f(x)

OpenStudy (gorv):

f(x) is the value of y which is multiplied by 6

OpenStudy (gorv):

x unchanged as we havent done anythin with x

OpenStudy (gorv):

new value of y= 6*y y=6*(-3)=-18

OpenStudy (anonymous):

Great thanks!